OpenStudy (amistre64):

What is the parametric equations of the a line parallel to <4,1,5> containing the point P(6,-4,8)?

OpenStudy (amistre64):

easy medal...you can do it :)

OpenStudy (amistre64):

x = Px + at y = Py + bt z = Pz + ct <a,b,c> ; P(x,y,z) right?